Computech Corporation is expanding rapidly and currently needs to retain all of its earnings; hence, it does not pay dividends. However, investors expect Computech to begin paying dividends, beginning with a dividend of $1.75 coming 3 years from today. The dividend should grow rapidly - at a rate of 41% per year - during Years 4 and 5, but after Year 5, growth should be a constant 4% per year. If the required return on Computech is 15%, what is the value of the stock today? Do not round intermediate calculations. Round your answer to the nearest cent.
Answer #1

D3=1.75

D4=(1.75*1.41)=2.4675

D5=(2.4675*1.41)=3.479175

Value after year 5=(D5*Growth rate)/(Required rate-Growth rate)

=(3.479175*1.04)/(0.15-0.04)

=$32.8940182

Hence current value=Future dividend and value*Present value of discounting factor(rate%,time period)

=1.75/1.15^3+2.4675/1.15^4+3.479175/1.15^5+32.8940182/1.15^5

=$20.65(Approx).